### v1

# FedSlate: A Federated Deep Reinforcement Learning Recommender System*

Abstract
Reinforcement learning methods have been used to optimize long-term user engagement in recommendation systems. However, existing reinforcement learning-based recommendation systems do not fully exploit the relevance of individual user behavior across different platforms. One potential solution is to aggregate data from various platforms in a centralized location and use the aggregated data for training. However, this approach raises economic and legal concerns, including increased communication costs and potential threats to user privacy. To address these challenges, we propose FedSlate, a federated reinforcement learning recommendation algorithm that effectively utilizes information that is prohibited from being shared at a legal level. We employ the SlateQ algorithm to assist FedSlate in learning users’ long-term behavior and evaluating the value of recommended content. We extend the existing application scope of recommendation systems from single-user single-platform to single-user multi-platform and address cross-platform learning challenges by introducing federated learning. We use RecSim to construct a simulation environment for evaluating FedSlate and compare its performance with state-of-the-art benchmark recommendation models. Experimental results demonstrate the superior effects of FedSlate over baseline methods in various environmental settings, and FedSlate facilitates the learning of recommendation strategies in scenarios where baseline methods are completely inapplicable. Code is available at https://github.com/TianYaDY/FedSlate.
Index Terms:
Recommender system, reinforcement learning, federated learning, vertical federated learning, privacy preservation.
I Introduction
Given the significant influence of users’ instant reactions to recommended content on their future behavior, research in the realm of content recommendation systems has increasingly adopted deep reinforcement learning (DRL) strategies. These strategies aim to optimize the balance between immediate user engagement and long-term retention [1, 2, 3]. Additionally, recent advancements in advertising recommendation systems have integrated reinforcement learning to achieve a balance between generating ad revenue and minimizing negative user experiences [4]. However, these systems often overlook an essential aspect: user behavior is not isolated but affected by recommendations from various platforms, indicating interdependencies between a user’s activities across different services. An obvious solution to leverage these behavioral correlations is to consolidate user data from multiple sources for the development of a unified model. Nevertheless, the introduction of stringent data privacy regulations, such as the General Data Protection Regulation (GDPR) in the European Union, the Personal Data Protection Act (PDPA) in Singapore, and the California Consumer Privacy Act (CCPA) in the United States, raises significant legal challenges. Directly employing user data could infringe upon privacy rights [5, 6]. Additionally, the communication overhead [7, 8, 9] poses a barrier to the straightforward implementation of centralized learning approaches. To circumvent these obstacles while harnessing the coherence in user behavior across platforms, we advocate for the adoption of federated learning (FL) techniques [10] to refine reinforcement learning-based recommendation algorithms.
In numerous contexts, the correlation between user behaviors yields substantial benefits. Within the financial sector, for instance, customers often engage with both stock trading and online payment services. A collaborative model trained by these services can effectively ascertain a customer’s risk profile, investment patterns, and spending behaviors, thereby facilitating tailored financial product recommendations that meet individual needs. Similarly, in the realm of online advertising, user interactions with multiple platforms reveal interconnected behaviors. A user may explore health foods on a social network while simultaneously shopping for wellness products on an e-commerce site. Utilizing FL, platforms can collectively develop a model that captures the user’s interests and buying inclinations, enhancing the precision of targeted advertising. Consequently, a recommendation system underpinned by FL presents an approach that optimizes privacy, minimizes communication overhead, and improves long-term value for users.
Recommendation systems commonly adopt a slate-based approach, wherein multiple items are simultaneously suggested to the user, allowing them to select and view their preferred item. This presents a significant challenge for the direct application of reinforcement learning (RL) due to the large action space involved. SlateQ [1] is a recommendation algorithm that utilizes the slate decomposition technique to address the challenge of recommending multiple items, known as a recommendation slate, to users. It effectively resolves the issue of the large action space faced by previous RL recommendation algorithms. However, SlateQ can only be deployed separately on different platforms and cannot be easily extended to the scenario of joint deployment across multiple platforms. As a result, it fails to leverage the correlations between user behaviors on different platforms. To address this problem, we propose FedSlate, assuming that a user’s behavior and response on one platform can be influenced by the recommended content from other platforms. Furthermore, certain influences are regarded as “inaccessible” to specific platforms. Specifically, we assume that certain agents are unable to directly receive rewards (even though the rewards exist), and these agents cannot make decisions based solely on their own information.
Our FedSlate algorithm incorporates both local and global models. However, unlike the adaptive personalized federated learning (APFL) algorithm [11], we do not blend the local and global models proportionally. Instead, we adopt a method similar to “Q-value sharing” [12], where the local models generate Q-values that are passed as inputs to the central server. The central server then computes the global Q-values used for content recommendation selection. FedSlate can be divided into the following stages. First, dedicated agents on each platform calculate local Q-values based on observed user states and candidate recommendation content states. They transmit these values to the central server. Next, the central server collects the received local Q-values as inputs to the global Q-network and calculates the corresponding global Q-values for each local agent. Finally, the central server distributes the Q-values to the respective agents, and the local agents make policy selections based on the received Q-values. It is important to note that each agent is unaware of the Q-network parameters of other agents. The central server calculates the global Q-values as many times as there are local agents since each agent has its own distinct “local Q-values”. These stages do not include the process of updating the local and global Q-networks. In FedSlate, we iteratively update the global network and local network based on the outputs of the global Q-network and the rewards obtained by an agent in its environment. During this process, the global network is updated multiple times, while the local network is updated only once.
We summarize the main contributions into threefold:
  1. 1.
    We introduce FL to address the issue of interdependence in user behavior due to the influence of recommendations from different platforms in content recommendation systems. By harnessing user behavior data for collective training from multiple sources, we enhance the performance of existing reinforcement learning-based recommendation algorithms. Our approach strategically balances privacy concerns with reduced communication overhead.
  2. 2.
    Prior research on FL in recommendation systems mainly focused on applying this technique to conventional algorithms, as evidenced by [13, 14, 15, 16]. These methods typically average local model outputs without accounting for user heterogeneity [17], which can be problematic given the diverse and imbalanced nature of user preferences. Such systems are often inadequate in reflecting long-term user interests. To overcome these limitations, we introduce FedSlate, a novel algorithm that marries the concepts of SlateQ and FL. FedSlate is adept at tracking long-term user behavior and assimilating the impact of cross-platform recommendations on user activity within a singular platform. By leveraging both local and global models and adopting a “Q-value sharing” technique, FedSlate facilitates the central server in generating comprehensive Q-values for optimal content recommendation.
  3. 3.
    We propose an innovative update mechanism specific to FL augmented with reinforcement learning in our FedSlate algorithm. Unlike existing adaptive personalization strategies in FL, FedSlate utilizes a unique sequential updating process. The global network is updated multiple times based on the global Q-network’s outputs, while the local networks receive a singular update following the local agent’s environment rewards. This approach significantly enhances the algorithm’s efficiency and output.
  4. 4.
    The proposed FedSlate framework is skilled at deriving recommendation strategies from data inaccessible to local agents, an attribute of significant relevance in practical implementations. This aspect is crucial in contexts where privacy issues or data governance policies restrict information access.
In Section II, we review relevant prior research related to our FedSlate algorithm. Subsequently, in Section III, we present the problems that FedSlate aims to resolve. Section IV describes the intricacies of the FedSlate algorithm. Finally, in Section V, we deploy RecSim [18] to establish a simulation environment for evaluating recommender systems and to examine the efficacy of the FedSlate algorithm in this context.
II Related Work
Recommender systems are a critical type of information filtering system that leverages user preference and behavior analysis to provide personalized suggestions [19]. These systems are extensively applied in various sectors, including e-commerce, social media, and entertainment platforms like music and video streaming, aiming to enhance content discovery and improve the overall user experience. Supervised learning (SL) techniques are commonly utilized in these systems to perform predictive and recommendatory functions, relying on patterns and rules derived from labeled training data. In this realm, training datasets, which include users’ historical interactions and their corresponding feedback or ratings, are instrumental. Among the prevalent SL-based methodologies for recommender systems, collaborative filtering (CF) stands out [20, 21]. User-based CF [22] suggests items by identifying similarities between users’ past behaviors, positing that users with comparable preferences are likely to be interested in similar items. Conversely, item-based CF [23], recommends based on item similarities. These CF methods are favored for their simplicity and proven effectiveness. Content-based recommendations [24] represent another widespread SL approach, employing item characteristics and user preferences to formulate suggestions. For example, a movie recommendation system may use a content-based method to recommend movies by considering aspects such as genre, actors, and directors, thus predicting a user’s potential interests. Furthermore, SL-based recommender systems may integrate various machine learning models, including decision trees [25], support vector machines (SVM) [26], and deep neural networks (DNN) [27]. These models enhance the recommendation process by adapting to diverse dataset features and characteristics while learning from users’ preferences and behavior patterns during training.
SL-based recommender systems have shown proficiency in short-term prediction tasks; however, incorporating reinforcement learning (RL) has been identified as critical for long-term prediction challenges [28, 29, 30]. RL, a machine learning paradigm, seeks to establish optimal behavioral policies through environmental interactions [31, 32] and distinguishes itself by concentrating on goal-directed decision-making, employing a trial-and-error process with a rewards system. This method excels in scenarios requiring foresight, such as financial investment [33] and complex planning, due to its adeptness at managing delayed rewards. In the medical domain, where uncertainty and dynamic conditions are prevalent, like ventilator management [34, 35], RL’s attributes prove exceptionally beneficial. Within RL-based recommendations, there are “model-based” and “model-free” methods; our focus is on the “model-free” category, which is straightforward to implement and yields superior long-term performance. Before this approach, DRN [36] implemented a deep Q-network (DQN) to create user profiles and an activity score. Subsequently, the social attentive deep Q-network (SADQN) [37] enhanced DQN with an attention mechanism to leverage social influences. Diverging from these, our contribution, FedSlate, leverages cross-platform user performance similarities. While some studies have adopted policy gradient methods, such as the Monte Carlo-based REINFORCE algorithm for large-scale recommendation environments [38], the SlateQ algorithm [1] is particularly influential in our approach. It decomposes the Q-value of a recommendation slate into individual item Q-values, effectively managing extensive action spaces and offering three item-wise Q-value-based selection strategies. Notwithstanding, existing RL-based systems primarily optimize for single-platform performance, neglecting the multiplatform influences on real-world users. To address this gap, we propose the incorporation of FL paradigms into RL-based recommendation systems.
FL involves training statistical models directly on devices to develop a joint model capable of generating data across distributed nodes [10]. Traditional distributed learning methods generally presuppose that local data samples are independently and identically distributed (IID); however, FL typically operates under the premise that data among clients is non-IID [39, 40]. Prior research has primarily utilized RL to enhance FL’s performance [41]. For instance, [41] employs a Deep Q-Learning (DQL) strategy to select devices for participation in successive communication rounds, thus minimizing the number of required rounds. Related work can be found in [42, 43]. In contrast, [44] applies a deep reinforcement learning (DRL) technique to modulate the CPU frequency of faster devices within an FL training cohort, balancing energy efficiency with training velocity. Distinct from these studies, our FedSlate algorithm “federates” DRL rather than simply applying RL to optimize FL. Concerning federated recommender systems, current studies [13, 14, 15, 16], have mainly adapted FL for traditional recommendation frameworks, setting them apart from our DRL-centric FedSlate approach. The most significant influence on our work is Federated deep Reinforcement Learning (FedRL) [12], which introduces an innovative DRL paradigm to collaboratively construct high-quality models for agents. Our approach adopts the “Q-value sharing” concept from FedRL to monitor long-term user behaviors and evaluate the impact of recommended content from various platforms on a user’s actions within a specific platform, thus tackling the complexities of applying RL recommendation algorithms in a multi-platform context. While FedRL advocates for a decentralized scheme, our FedSlate algorithm is designed with centralization in mind but can also be adapted to a decentralized format for practical applications.
III Problem Definition
In this section, we introduce an augmented Markov Decision Process (MDP) model tailored for the single-user, multi-platform context. This model captures the dynamics wherein platforms employ recommendation systems to curate slates of content. Users engage with these slates by selecting an item—or opting out—and post-consumption, decide whether to seek additional recommendations or end their session. It is important to note that users may transition across platforms in pursuing content that piques their interest, a pattern that closely reflects real-world user behavior. User responses to content are multifaceted, encompassing metrics such as browsing duration, “likes”, and comments. However, for the sake of a generalized model, we limit our focus to user engagement as the singular metric of reward. Subsequently, we outline the assumptions underpinning our problem, some aligning with the SlateQ framework [1] and others specific to the single-user, multi-platform scenario. We conclude this section by detailing a precise formalization of our federated reinforcement learning recommendation problem.
III-A An Extended MDP Model for Slate Recommendation
The recommendation and user interaction behaviors within a single platform are aptly modeled by an MDP, characterized by states $S$, actions $A$, a reward function $R$, a transition kernel $P$, and a discount factor $\gamma$ [1]. We will now elucidate the critical elements of this model:
  • The state $\mathcal{S}$ implements the user’s condition, comprising both observable attributes such as age, gender, and self-reported interests, and historical interactions including prior browsing activity and responses to earlier recommendations.
  • Action $\mathcal{A}$ encompasses all potential recommendation arrays; upon generating a collection of $\mathcal{I}$ items, the system is tasked with curating a subset of $k$ items to form the user’s slate, denoted by ${A\subseteq\mathcal{I}\quad s.t.\left|A\right|=k}$, where $k$ is the pre-defined slate size.
  • The transition probability $P(s{}^{\prime}|s,A)$ quantifies the likelihood of migrating from state $s$ to state $s{}^{\prime}$ subsequent to action $A$.
  • The reward $R(s,A)$ assesses the anticipated user engagement with the slate $A$, serving as an index of the user’s interaction with the recommended items.
The value function or Q-function of the policy $\pi:S\to A$ that the agent takes after observing state $s$ is given by the following equation:
$V^{\pi}\left(s\right)=R\left(s,\pi\left(s\right)\right)+\gamma\sum_{s{}^{\prime}\in S}P\left(s{}^{\prime}|s,\pi\left(s\right)\right)V^{\pi}\left(s{}^{\prime}\right)$ (1)
$Q^{\pi}\left(s,A\right)=R\left(s,A\right)+\gamma\sum_{s{}^{\prime}\in S}P\left(s{}^{\prime}|s,A\right)V^{\pi}\left(s{}^{\prime}\right)$ (2)
The optimal policy $\pi^{\ast}$ maximizes the expected value $V\left(s\right)$. Therefore, we focus on the following expression:
$V^{\ast}\left(s\right)=\mathop{max}_{A\in\mathcal{A}}R\left(s,A\right)+\gamma\sum_{s{}^{\prime}\in S}P\left(s{}^{\prime}|s,A\right)V^{\ast}\left(s{}^{\prime}\right)$ (3)
$Q^{\ast}\left(s,A\right)=R\left(s,A\right)+\gamma\sum_{s{}^{\prime}\in S}P\left(s{}^{\prime}|s,A\right)V^{\ast}\left(s{}^{\prime}\right)$ (4)
Within this framework, the optimal policy $\pi^{\ast}$ is defined such that $\pi^{\ast}(s)=\underset{A\in\mathcal{A}}{\mathrm{argmax}},Q^{\ast}(s,A)$.
We devised a user case where an individual user navigates two distinct platforms to demonstrate our algorithm, a scenario that is scalable to more platforms. However, the MDP model as initially conceived does not accommodate a user engaging with multiple platforms. To address this limitation, we propose an augmented model involving platforms A and B, wherein the MDP framework is applied independently to each. Concretely, this involves discrete recommendation agents for each platform (agent $\alpha$ for platform A and agent $\beta$ for platform B), with corresponding user states ($S_{\alpha}$ and $S_{\beta}$), agent actions ($A_{\alpha}$ and $A_{\beta}$), transition probabilities ($P_{\alpha}(s^{\prime}|s,A)$ and $P_{\beta}(s^{\prime}|s,A)$), rewards ($R_{\alpha}$ and $R_{\beta}$), and Q-functions ($Q_{\alpha}$ and $Q_{\beta}$).
III-B Necessary Assumptions
In our federated reinforcement learning recommendation problem, we make the following assumptions:
  • A1: A user selects only one item at a time (or may choose not to select, represented as $\bot$ for a null item).
  • A2: Transitions depend solely on the selection. Specifically, the user’s state changes, and a reward (user engagement) is generated only when the user consumes item $i$. Additionally, while a user engages with a platform, the states of other platforms remain frozen.
  • A3: There is interconnectedness between the user’s behaviors on different platforms, and the impact of a single platform on the user is “cross-platform”.
  • A4: After consuming item $i$, the user’s interest transfers to other platforms.
  • A5: Only the output values of $Q_{\alpha}$ and $Q_{\beta}$ are shared for learning the joint policy $\pi^{\ast}_{fed}$. Other information, including $D_{\alpha}$ and $D_{\beta}$, is locally visible only.
A1 is the original assumption in SlateQ (SC). A2 extends the RTDS assumption in SlateQ to the multi-platform scenario. A3, A4, and A5 are specific assumptions for the single-user multi-platform context. A3 is the most important premise of our method, which is intuitive and easily acceptable. Note that A4 is a simplification of actual user behavior since regular users often switch platforms after consuming a “series” of items. However, we believe this simplification brings about concise algorithmic descriptions, and we consider A4 to be easily relaxed (by fixing the states of platforms where the user is not present and allowing state transitions to continue on the user’s current platform). A5 ensures that information about the user on different platforms is not leaked (we believe that even for the same user, information should not be shared between platforms without user authorization).
III-C Recommendation Problem
After extending the original MDP model, we can formally define our recommendation problem based on A1-A5. The existing platforms $\alpha$ and $\beta$ take turns randomly recommending slates to customers and recording transitions. This results in a series of transitions $D_{\alpha}=\left\{\left\langle s_{\alpha},A_{\alpha},s{}^{\prime}_{\alpha},r_{\alpha}\right\rangle\right\}$ for agent $\alpha$ and transitions $D_{\beta}=\left\{\left\langle s_{\beta},A_{\beta}\right\rangle\right\}$ for agent $\beta$, where $D_{\alpha}$ and $D_{\beta}$ are one-to-one correspondence. Our goal is to learn a joint policy $\pi^{\ast}$ that, based on $s_{\alpha}$ and $s_{\beta}$, maximizes the lifetime value (LTV) across all platforms. It should be noted that platform B does not record the user’s response, i.e., $D_{\beta}$ does not contain $r_{\beta}$, which deviates from the previous MDP model. We adopt this setting because certain platforms may not have direct access to user feedback on recommended content (although the platform’s impact on the user is real). User preferences or aversions may be reflected on other platforms. We aim to demonstrate the friendliness of FedSlate towards these “unavailable feedback” platforms—even if a platform cannot directly update its recommendation strategy based on user feedback, it can still benefit from performance improvements in the federation.
IV Our FedSlate Method
In this section, we will provide a detailed description of our FedSlate method. We utilize the SlateQ algorithm to assist us in evaluating the value of recommended content and tracking user feedback over the long term. Therefore, we will begin by briefly introducing the original SlateQ algorithm. Subsequently, we will present the components of our algorithm, followed by a description of the specific details of the algorithm. Lastly, we will propose an extended version of our algorithm to address situations where team rewards are excessively sparse.
IV-A SlateQ Algorithm
In Section 3.1, we presented the MDP model for the recommendation problem. In the case of a single user and a single platform MDP, the original SlateQ algorithm aims to find an optimal policy $\pi^{\ast}$ that satisfies $\pi^{\ast}\left(s\right)=argmax_{A\in\mathcal{A}}Q^{\ast}\left(s,A\right)$. However, in Slate recommendation problems, the action space becomes extremely large, resulting in excessive computational and time resource requirements. Specifically, if we attempt to select k items from a set of $\mathcal{I}$ items to form a slate and consider the impact of the position of recommended items within the slate on user feedback, the action space would be of size ${A_{\mathcal{I}}^{k}}$.
To address this issue, SlateQ decomposes $Q^{\pi}\left(s,A\right)$ and represents slate-level Q-values as item-level Q-values $\overline{Q}^{\pi}\left(s,i\right)$, significantly reducing the agent’s action space. The decomposition is achieved using the following formula:
$Q^{\pi}\left(s,A\right)=\sum_{i\in A}P\left(i|s,A\right)\overline{Q}^{\pi}\left(s,i\right)$ (5)
The decomposed Q-values can be updated using a simple Temporal Difference (TD) method:
$\overline{Q}^{\pi}(s,i)\leftarrow\alpha(r+\gamma\sum_{j\in A^{\prime}}P(j|s^{\prime},A^{\prime})\overline{Q}^{\pi}(s^{\prime},j))+(1-\alpha)\overline{Q}^{\pi}\left(s,i\right)$ (6)
To fully satisfy the requirements of Q-learning, it is only necessary to introduce the usual maximization step:
$\overline{Q}(s,i)\leftarrow\alpha(r+\gamma\mathop{max}_{A^{\prime}\in\mathcal{A}}\sum_{j\in A^{\prime}}P(j|s^{\prime},A^{\prime})\overline{Q}(s^{\prime},j))+(1-\alpha)\overline{Q}\left(s,i\right)$ (7)
Please note that SlateQ assumes the user choice model $P(i|s,A)$ is known. Models such as MNL, CL, and cascade can easily be learned using user response data, and this does not depend on LTV.
SlateQ offers multiple strategies to select recommended slates based on item Q-values. We consider the trade-off between computational and time resources and the effectiveness of the strategy, and we will only provide a detailed explanation of the Greedy optimization approach. A simple approach to utilizing item Q-values in slate construction is to use the Q-values as item scores, sort the items in descending order of scores, and select the top $k$ items to form the slate. However, this approach, known as the Top-$k$ method, fails to capture the influence of the first $L-1$ items on the $L$th slot (for $1<L\leq k$). Greedy optimization differs from the aforementioned method as it updates the item scores based on the current partial slate. For example, given $A^{\prime}=\{i_{(1)},...,i_{(L-1)}\}$ of size $L-1<k$, the $L$th item is selected based on the maximum marginal value it provides:
$\mathop{argmax}_{i\notin A^{\prime}}\frac{v(s,i)\overline{Q}(s,i)+\textstyle\sum_{l<L}v(s,i_{(l)})\overline{Q}(s,i_{(l)})}{v(s,i)+v(s,\perp)+\textstyle\sum_{l<L}v(s,i_{(l)})}$ (8)
IV-B The FedSlate Algorithm
In this section, we will introduce our FedSlate algorithm in a bottom-up manner, starting from some necessary components.
Basic Q Networks
We establish two Q networks, denoted as $Q_{\alpha}\left(s_{\alpha};{\theta}_{\alpha}\right)$ and $Q_{\beta}\left(s_{\beta};{\theta}_{\beta}\right)$, for agents $\alpha$ and $\beta$ respectively. Here, ${\theta}_{\alpha}$ and ${\theta}_{\beta}$ represent the parameters of the Q networks. It should be noted that both Q networks have the same structure as the Q network in the original SlateQ algorithm. Their output is a tensor of the same size as the number of candidate documents $\mathcal{I}$ (equivalent to the set $\overline{Q}(s,i),i\in\mathcal{I}$). However, we do not directly use their output values for slate recommendations. Instead, we use the output values as inputs for the federated Q network.
Federated Agent
To exploit information from both Platform A and Platform B, we introduce a third agent, referred to as agent $fed$, which receives the output values of $Q_{\alpha}$ and $Q_{\beta}$.
Within agent $fed$, we construct a simple multi-layer perceptron (MLP) module, also known as the federated Q network mentioned earlier, denoted as $Q^{f}$. This network utilizes the output of the two basic Q networks to derive Q values specifically used for slate selection. When $Q^{f}$ is employed for slate content selection, it differs depending on whether it is used by agent $\alpha$ or agent $\beta$. Specifically, agent $\alpha$ and agent $\beta$ have their own respective output values from $Q^{f}$, denoted as $Q_{\alpha}^{f}$ and $Q_{\beta}^{f}$, defined as follows:
$Q_{\alpha}^{f}(\cdot;\theta_{\alpha},\theta_{\beta},\theta_{f})=MLP([Q_{\alpha}(s_{\alpha};\theta_{\alpha})|Q_{\beta}(s_{\beta};\theta_{\beta});\theta_{f})$ (9)
$Q_{\beta}^{f}(\cdot;\theta_{\alpha},\theta_{\beta},\theta_{f})=MLP([Q_{\beta}(s_{\beta};\theta_{\beta})|Q_{\alpha}(s_{\alpha};\theta_{\alpha});\theta_{f})$ (10)
Where $\theta_{f}$ represents the parameters of the MLP, and [|][$\cdot$|$\cdot$][ ⋅ | ⋅ ] denotes the concatenation operation. In the above equation, we utilize the first position of the MLP input to represent “one’s own Q value”, while the second position represents “Q values that do not belong to oneself”.
During the actual process of slate recommendation and Q network update, we fix the parameters $\theta$ of the Q network for agents on the platform where the user is not present. We treat the output of the Q network as a constant to ensure the stability of our algorithm during the learning phase,
$Q_{\alpha}^{f}(\cdot,C_{\beta};\theta_{\alpha},\theta_{f})=MLP([Q_{\alpha}(s_{\alpha};\theta_{\alpha})|C_{\beta});\theta_{f})$ (11)
$Q_{\beta}^{f}(\cdot,C_{\alpha};\theta_{\beta},\theta_{f})=MLP([Q_{\beta}(s_{\beta};\theta_{\beta})|C_{\alpha});\theta_{f})$ (12)
Where $C_{\alpha}=Q_{\alpha}(s_{\alpha};\theta_{\alpha})$ and $C_{\beta}=Q_{\beta}(s_{\beta};\theta_{\beta})$ are the fixed outputs of the basic Q networks that we mentioned earlier [^1] Please note that, for the sake of brevity in expression, we will not differentiate between the Q networks themselves and their output values in the following text. We will uniformly use $Q$ to represent them..
Agent $fed$ is responsible for training the Q-networks. During the training phase, agent $fed$ sequentially receives $Q_{\alpha}$ and $Q_{\beta}$ and updates the corresponding networks. In order to minimize the error, we use the Huber loss to define the loss functions $L_{\alpha}(\theta_{\alpha},\theta_{f})$ and $L_{\beta}(\theta_{\beta},\theta_{f})$ for agents $\alpha$ and $\beta$ respectively:
$\mathcal{L}_{\delta}(a)=\left\{\begin{array}[]{ll}\frac{1}{2}a^{2}&\text{ for }|a|\leq\delta,\\ \delta\cdot\left(|a|-\frac{1}{2}\delta\right),&\text{ otherwise. }\end{array}\right.$ (13)
$L_{\alpha}(\theta_{\alpha},\theta_{f})=\frac{1}{|B|}\sum\mathcal{L}(Y_{\alpha}-Q_{f}^{\alpha}(\cdot,C_{\beta};\theta_{\alpha},\theta_{f}))$ (14)
$L_{\beta}(\theta_{\beta},\theta_{f})=\frac{1}{|B|}\sum\mathcal{L}(Y_{\alpha}-Q_{f}^{\beta}(\cdot,C_{\alpha};\theta_{\beta},\theta_{f}))$ (15)
Where $|B|$ represents the batch size during training, $Y_{\alpha}=r_{\alpha}+\gamma\mathop{max}\limits_{A_{\alpha}^{\prime}\in\mathcal{A_{\alpha}}}\sum_{j\in A_{\alpha}^{\prime}}P(j|s_{\alpha}^{\prime},A_{\alpha}^{\prime})Q_{\alpha}^{f}(s_{\alpha}^{\prime},j)$ denotes the target Q-value. It is crucial to note that in Eq.(15), the updates of $Q_{\alpha}$ and $Q_{\beta}$ are both contingent upon $r_{\alpha}$ due to agent $\beta$ lacking access to $r_{\beta}$.
Overview of Acting and Learning
Our FedSlate algorithm can be divided into two parts: “acting” and “learning”, as shown in Fig.1. In the loop of our algorithm, we first perform several rounds of “acting” to recommend slates to users and record their feedback. Then, we execute one iteration of “learning” to update the network using the collected experiences, thereby optimizing the recommendation strategy. This loop continues for multiple iterations during the training phase. Due to the modular design of our algorithm, we do not strictly differentiate between the training and testing phases. When there is no need for policy optimization, we simply skip the execution of the “learning” module in the loop. Similarly, if there are changes in the distribution of input data for the recommendation algorithm (e.g., business adjustments on the platform), the “learning” module can be reactivated.
Figure 1: Main Loop Process of the FedSlate Algorithm.
As shown in Fig.2,in the “acting” phase, agent $fed$ initiates inquiry requests to agent $\alpha$ and $\beta$. Agent $\alpha$ and $\beta$ calculate $Q_{\alpha}$ and $Q_{\beta}$ based on their current states $s_{\alpha}$ and $s_{\beta}$, respectively, and send them to agent $fed$. Agent $fed$ computes $Q_{\alpha}^{f}$ and $Q_{\beta}^{f}$ and sends them back to agents $\alpha$ and $\beta$, respectively. Agent $\alpha$ and $\beta$ construct slates using a greedy method based on the received Q-values and recommend them to the users.
During the “learning” phase, agents $\alpha$ and $\beta$ are assigned random indices $IDs$ of size $|B|$ by agent $fed$, which correspond to specific training batches ($D_{\alpha}$ and $D_{\beta}$ are in a one-to-one correspondence). Utilizing these batches, agents $\alpha$ and $\beta$ calculate $Q_{\alpha}$, $Q_{\alpha}^{\prime}$, and $Q_{\beta}$, and subsequently transmit these values to agent $fed$ for the computation of $Q_{f}^{\alpha}$ and ${Q_{f}^{\alpha}}^{\prime}$. Notably, $Q_{\alpha}^{\prime}$ and ${Q_{f}^{\alpha}}^{\prime}$ denote the subsequent Q-values. Agent $fed$ then returns $Q_{f}^{\alpha}$ and ${Q_{f}^{\alpha}}^{\prime}$ to agent $\alpha$ for the derivation of $Y_{\alpha}$ and the updating of the networks $Q_{\alpha}(s_{\alpha};\theta_{\alpha})$ and $Q_{f}$. After these updates, $Y_{\alpha}$ is conveyed to agent $\beta$, while agent $\alpha$ formulates a refreshed $Q_{\alpha}$ using the newly updated network to forward to agent $fed$. Utilizing the updated $Q_{\alpha}$ and the initial $Q_{\beta}$, agent $fed$ calculates $Q_{\beta}^{f}$ and dispatches it to agent $\beta$, who then updates the networks $Q_{\beta}(s_{\beta};\theta_{\beta})$ and $Q_{f}$ with the aid of $Y_{\alpha}$ and $Q_{\beta}^{f}$. The learning protocol mandates a single update for each local network and two updates for the global network. Figure 3 provides a schematic representation of the “learning” process.
Figure 2: Overview of the Acting Component in the FedSlate Algorithm.
The detailed acting and learning processes can be found in Algorithm 1, 2 and 3, where we will introduce some crucial details. Firstly, in step 21 of Algorithm 1, the agent $\alpha$ computes $Y_{\alpha}$ using Eq.(7),(8). The inputs $Q_{f}^{\alpha}$ and ${Q_{f}^{\alpha}}^{\prime}$ received by agent $\alpha$ are both tensors of size $[B,\mathcal{I}]$, where $\mathcal{I}$ represents the size of the candidate documents. Secondly, since we employ the Q-values of items to update our Q-network, a certain transformation must be applied to $Q_{f}^{\alpha}$ and ${Q_{f}^{\alpha}}^{\prime}$. We only consider the Q-values of items that have been actually selected by the user as the online Q-values (the left-hand side of Eq.(7)), while disregarding the Q-values of items that were not chosen by the user. Thirdly, in order to calculate the target Q-values (the right-hand side of Eq.(7)), we first generate a slate, then compute the probabilities for each recommended item on the slate, and finally take the inner product of these probabilities with their corresponding item’s Q-values. This inner product serves as the target Q-value and is utilized in step 21 of Algorithm 2 to address the problem of agent $\beta$ being unable to determine the online Q-values for TD updates based on the user’s selected item, given the assumption that agent $\beta$ cannot access user feedback.
Figure 3: Overview of the Learning Component in the FedSlate Algorithm.
Algorithm 1 FedSlate-ALPHA
0:  $S_{\alpha}$
0:  $None$
1:  function $Init()$
2:      Initialize $Q_{\alpha}$ with random values for $\theta_{\alpha}$
3:  end function
4:  function $ComputeQAlpha()$
5:      Observe $s_{\alpha}$
6:      Compute $Q_{\alpha}(s_{\alpha};\theta_{\alpha})$
7:      return $Q_{\alpha}$
8:  end function
9:  function $RecommendSlate(Q_{\alpha}^{f})$
10:      Observe $\mathcal{I}_{\alpha}$
11:      Construct Slate $A_{\alpha}$ based on Eq.(8)
12:      Recommend Slate $A_{\alpha}$, obtain state $s_{\alpha}^{\prime}$ and reward $r_{\alpha}$
13:      Store $(s_{\alpha},A_{\alpha},r_{\alpha},s_{\alpha}^{\prime})$ in $D_{\alpha}$
14:  end function
15:  function $ComputeQAlphaBatch(IDs)$
16:      Sample batch of $D_{\alpha}$ based on indices $IDs$
17:      Compute $Q_{\alpha}(s_{\alpha};\theta_{\alpha})$ and $Q_{\alpha}^{\prime}(s_{\alpha}^{\prime};\theta_{\alpha}^{\prime})$
18:      return batches of $Q_{\alpha}$, $Q_{\alpha}^{\prime}$
19:  end function
20:  function $UpdateQNet(Q_{f}^{\alpha},{Q_{f}^{\alpha}}^{\prime})$
21:      Compute $Y_{\alpha}$ based on Eq.(7,8)
22:      Update $Q_{\alpha}$ and $Q_{f}$ based on Eq.(7)
23:      return $Y_{\alpha}$
24:  end function
Algorithm 2 FedSlate-BETA
0:  $S_{\beta}$
0:  $None$
1:  function $Init()$
2:      Initialize $Q_{\beta}$ with random values for $\theta_{\beta}$
3:  end function
4:  function $ComputeQBeta()$
5:      Observe $s_{\beta}$
6:      Compute $Q_{\beta}(s_{\beta};\theta_{\beta})$
7:      return $Q_{\beta}$
8:  end function
9:  function $RecommendSlate(Q_{\beta}^{f})$
10:      Observe $\mathcal{I}_{\beta}$
11:      Construct Slate $A_{\beta}$ based on Eq.(8)
12:      Recommend Slate $A_{\beta}$
13:      Store $(s_{\beta},A_{\beta})$ in $D_{\beta}$
14:  end function
15:  function $ComputeQBetaBatch(IDs)$
16:      Sample batch of $D_{\beta}$ based on indices $IDs$
17:      Compute $Q_{\beta}(s_{\beta};\theta_{\beta})$
18:      return batches of $Q_{\beta}$
19:  end function
20:  function $UpdateQNet(Q_{f}^{\beta},Y_{\alpha})$
21:      Update $Q_{\beta}$ and $Q_{f}$ based on Eq.(7)
22:  end function
Algorithm 3 FedSlate-FED
0:  Boolean value to determine whether to cancel the learn module: $is\_learn$,Integer type representing the learning interval: $learn\_every$
0:  $None$
1:  Initialize $Q_{f}$ with random values for $\theta_{f}$
2:  Call FedSlate-ALPHA.$Init()$,FedSlate-BETA.$Init()$
3:  for episode = $1$ to $M$ do
4:     Initialize $step=0$
5:     while True do
6:        Call $Q_{\alpha}=$ FedSlate-ALPHA.$ComputeQAlpha()$
7:        Call $Q_{\beta}=$ FedSlate-BETA.$ComputeQBeta()$
8:        Compute $Q_{f}^{\alpha}$, $Q_{f}^{\beta}$ according to Eq.(11,12)
9:        Call FedSlate-ALPHA.$RecommendSlate(Q_{\alpha}^{f})$
10:        Call FedSlate-BETA.$RecommendSlate(Q_{\beta}^{f})$
11:        if $is\_learn=True$, $step\bmod learn\_every=0$ then
12:           Generate $IDs$ randomly
13:           Call $Q_{\alpha}$, $Q_{\alpha}^{\prime}$ $=$ FedSlate-ALPHA.$ComputeQAlphaBatch(IDs)$
14:           Call $Q_{\beta}$ $=$ FedSlate-BETA.$ComputeQBetaBatch(IDs)$
15:           Compute $Q_{\alpha}^{f}$, ${Q_{\alpha}^{f}}^{\prime}$ based on Eq.(11)
[…]
17:           Call $Q_{\alpha}$ $=$ FedSlate-ALPHA.$ComputeQAlphaBatch(IDs)$
18:           Compute $Q_{\beta}^{f}$ based on Eq.(12)
19:           Call FedSlate-BETA.$UpdateQNet(Q_{\beta}^{f},Y_{\alpha})$
20:        end if
21:        if terminal then
22:           break
23:        end if
24:        Let $step=step+1$
25:     end while
26:  end for
Algorithm 4 FedSlate-FED(extended)
0:  Boolean value to determine whether to cancel the learn module: $is\_learn$,Integer type representing the learning interval: $learn\_every$
0:  $None$
1:  Initialize $Q_{f}$ with random values for $\theta_{f}$
2:  Call FedSlate-ALPHA.$Init()$,FedSlate-BETA.$Init()$
3:  for episode=$1$ to $M$ do
4:     Initialize $step=0$
5:     while True do
6:        Call $Q_{\alpha}=$ FedSlate-ALPHA.$ComputeQAlpha()$
7:        Call $Q_{\beta}=$ FedSlate-BETA.$ComputeQBeta()$
8:        Compute $Q_{f}^{\alpha}$, $Q_{f}^{\beta}$ according to Eq.(11,12)
9:        Call FedSlate-ALPHA.$RecommendSlate(Q_{\alpha}^{f})$
10:        Call FedSlate-BETA.$RecommendSlate(Q_{\beta}^{f})$
11:        if $is\_learn=True$, $step\bmod learn\_every=0$ then
12:           Generate $IDs$ randomly
13:           Call $Q_{\alpha}$, $Q_{\alpha}^{\prime}$ $=$ FedSlate-ALPHA.$ComputeQAlphaBatch(IDs)$
14:           Call $Q_{\beta}$, $Q_{\beta}^{\prime}$ $=$ FedSlate-BETA.$ComputeQBetaBatch(IDs)$
15:           Compute $Q_{\alpha}^{f}$, ${Q_{\alpha}^{f}}^{\prime}$ based on Eq.(11)
[…]
17:           Call $Q_{\alpha}$ $=$ FedSlate-ALPHA.$ComputeQAlphaBatch(IDs)$
18:           Compute $Q_{\beta}^{f}$, ${Q_{\beta}^{f}}^{\prime}$ based on Eq.(12)
19:           Call FedSlate-BETA.$UpdateQNet(Q_{\beta}^{f},{Q_{\beta}^{f}}^{\prime})$
20:        end if
21:        if terminal then
22:           break
23:        end if
24:        Let $step=step+1$
25:     end while
26:  end for
In addition, as mentioned by [45], in real-world multi-agent systems, team rewards suffer from sparsity, making it difficult for algorithms to learn a successful team strategy to enhance overall reward. In our setting, if the rewards from Platform A and Platform B are too sparse (in other words, the information received by users on Platform A does not affect their reactions on Platform B), it is not possible to train multiple agents simultaneously using the reward from Platform A. Therefore, our algorithm has a simple variant to address this issue: Algorithm 1 is duplicated to replace Algorithm 2, and some minor modifications are made to Algorithm 3. Please refer to Algorithm 4 for the specific changes made to Algorithm 3.
By simple extension, our algorithm is capable of handling the sparsity issue of team rewards. However, it should be noted that the extended algorithm requires platform B to have access to rewards as well.
V Experimental Setup
In this section, the RecSim platform is utilized to construct a simulation environment aimed at evaluating the efficacy of our FedSlate algorithm. Specifically, detailed information regarding the simulation environment is provided. Subsequently, a comparative analysis is conducted between our algorithm and the SlateQ method, to ascertain whether agent $\alpha$ demonstrates enhanced performance subsequent to its participation in the FL process, as compared to its individual learning performance. A metric is proposed to assess this particular aspect. Furthermore, a comparison is made between FedSlate and a purely random recommendation approach, in order to ascertain if agent $\beta$ (as agent $\beta$ lacks access to user feedback and therefore cannot optimize the recommendation strategy using conventional methods) can derive benefits from the federated setting. By considering these two aspects, we aim to demonstrate whether our algorithm can effectively exploit the consistency of user behavior across different platforms. Lastly, to address the potential sparsity issue of team rewards, the performance of our extended FedSlate algorithm is evaluated within a more complex scenario and compared against the performance of the basic FedSlate algorithm under conditions of sparse team rewards.
V-A Simulation Environment
RecSim[18] is a simulation platform for constructing and evaluating recommendation systems that naturally support sequential interactions with users. Developed by Google, it simulates users and environments to assess the effectiveness and performance of recommendation algorithms. We employ RecSim to create an environment that reflects user behavior and item structure to evaluate our FedSlate algorithm.
We construct a “Choc vs. Kale” recommendation scenario, where the goal is to maximize user satisfaction and engagement over the long term by recommending a certain proportion of “chocolate” and “kale” elements. In this scenario, the “chocolate” element represents content that is interesting but not conducive to long-term satisfaction, while the “kale” element represents relatively less exciting but beneficial content for long-term satisfaction. The recommendation algorithm needs to balance these two elements to achieve maximized long-term user satisfaction. We believe this scenario aligns well with our assumption that “user responses to content on other platforms are influenced to some extent by the content they are exposed to on the current platform”. If a platform consistently recommends “chocolate” content to users, their long-term satisfaction is likely to be compromised.
In our scenario, the entire simulation environment consists primarily of document models and user models. The document model serves as the main interface for interaction between users and the recommendation system (agent) and is responsible for selecting a subset of documents from a database containing a large number of documents to deliver to the recommendation system. The user model simulates user behavior and reacts to the slates provided by the recommendation system.
The database in the document model essentially serves as a container for observable and unobservable features of underlying documents. In our scenario, document attributes are modeled as continuous features with values in the range of $[0,1]$, referred to as the Kaleness scale. A document with a score of 0 represents pure “chocolate”, which is interesting but regretful, while a document with a score of 1 represents pure “kale”, which is less exciting but nutritious. Additionally, each document has a unique integer ID, and the document model selects $N$ candidate documents in sequential order based on their IDs.
The user model includes unobservable and observable features of users. Based on these features, the model responds to the received slate according to certain rules. Each user is characterized by the features of net kale exposure ($nke_{t}$) and satisfaction ($sat_{t}$), which are associated through the sigmoid function $\sigma$ to ensure that $sat_{t}$ is constrained within a bounded range. Specifically, the satisfaction level is modeled as a sigmoid function of the net kale exposure, which determines the user’s satisfaction with the recommended slate:
$sat_{t}=\sigma(\tau\cdot nke_{t})$ (16)
Where, $\tau$ is a user-specific sensitivity parameter. Upon receiving a Slate from the recommendation system, users select items to consume based on the Kaleness scale of the documents. Specifically, for item $i$, the probability of it being chosen is determined by $p\sim e^{1-kaleness(i)}$. After making their selections, the net kale exposure evolves as follows:
$nke_{t+1}=\beta\cdot nke_{t}+2(k_{i}-1/2)+\mathcal{N}(0,\eta)$ (17)
Where, $\beta$ represents a user-specific memory discount, while $k_{i}$ corresponds to the kaleness of the selected item, and $\eta$ denotes some noise standard deviation. Lastly, our focus will be on the user’s engagement $s_{i}$, i.e. a log-normal distribution with parameters linearly interpolating between the pure kale response $(\mu_{k},\sigma_{k})$ and the pure choc response $(\mu_{c},\sigma_{c})$:
$s_{i}\sim log\mathcal{N}(k_{i}\mu_{k}+(1-k_{i})\mu_{c},k_{i}\sigma_{k}+(1-k_{i})\sigma_{c})$ (18)
The satisfaction variable $sat_{t}$ represents the sole dynamic component of the user’s state, and thus, we generate the user’s observable state based on it. In the simulation, user satisfaction is modeled and computed as a latent state. However, to simulate real-world scenarios, we map the latent state to an observable state by introducing noise to account for user uncertainty.
We will develop two distinct document models, representing Platform A and Platform B, respectively, and integrate them with a user model to establish a comprehensive environment. Furthermore, we will introduce a time budget parameter to the user model, constraining the browsing duration for users. Upon a user’s cessation of browsing, the environment will generate a terminal signal to indicate the completion of the interaction.
The evaluation environment for the baseline method (i.e., SlateQ) is intentionally simplified compared to the previously described environments. It comprises a single document model and a single user model, allowing for the simulation of sequential interaction behavior of an individual user on a solitary platform.
V-B Algorithm Evaluation
We evaluate our algorithm in a simulated environment,the schematic representation of the interactive process between the environment and the agent is depicted as illustrated in Fig.4. First, let’s define the states, actions, and rewards.
Figure 4: Interaction Process between Environment and Agent.
  • States: The environment state observed by Agent $\alpha$ consists of the user’s observable state (with a count of $1$) and the features of candidate documents (with a count of $N$). Since Platform A has access to user feedback information, we incorporate the user’s historical engagement records into the observable state. Specifically, we include the user’s previous 5 engages in the state. Therefore, the state $\alpha$ corresponding to Platform A is a tensor of size $[1+5\times n+N]$, where $n$ represents the slate size we set. Platform B, which lacks access to user feedback information, does not include the user’s historical engagement records in the observed environment state. Hence, the state $\beta$ corresponding to Platform B is a tensor of size $[1+N]$, containing only the current user’s observable state and the features of candidate documents.
  • Actions: The actions for both Agent $\alpha$ and Agent $\beta$ are similar. They involve recommending a slate of content determined by the algorithm. The actions are essentially an integer tensor of size $[n]$, representing the specific item IDs that form the slate.
  • Rewards: As our objective is to maximize long-term user satisfaction, we employ cumulative user engagement as the reward. It is important to note that the environment provides both user engages on Platform A and Platform B simultaneously. However, for Agent $\beta$, the engages on Platform B are completely inaccessible. We include this information in the environment solely for evaluating the performance of our algorithm and not for training our model.
  • Criteria (Agent $\alpha$): To assess the learning time consumption of Agent $\alpha$ in our algorithm, we introduce a metric called “Episodes to Reach Optimal Reward” (ETROR), denoted as $M^{{}^{\prime}}$. Let $M_{1}$ represent the number of episodes undergone by the baseline method (SlateQ) during training, and $M_{2}$ represent the number of episodes undergone by FedSlate during training. Similarly, we define $R_{M_{1}}$ and $R_{M_{2}}$ as the rewards achieved at time $M$ during the training process. For the baseline method, if $R_{M_{1}}\geq R_{M_{1}+t},t\in\mathbb{N}^{+}$, we consider $M_{1}=M_{1}^{{}^{\prime}}$. Taking into account the instability of reinforcement learning, we add a small positive integer term $\varepsilon$ to the inequality, i.e., $R_{M_{1}}+\varepsilon\geq R_{M_{1}+t},t\in\mathbb{N}^{+}$. In such cases, we consider $M_{1}$ as the number of episodes consumed by the baseline method to reach optimal rewards. Similar definitions can be applied to determine $M_{2}^{{}^{\prime}}$.
  • Criteria (Agent $\beta$): We denote $R_{M_{3}^{{}^{\prime}}}$ as the optimal reward achieved by Agent $\beta$ with the introduction of FL at $M_{3}^{{}^{\prime}}$. We define $R_{rnd}$ as the average rewards obtained by Platform B using a random recommendation method over episodes $[0,M_{3}^{{}^{\prime}}]$. If $R_{M_{3}^{{}^{\prime}}}\geq R_{rnd}$, we conclude that Agent $\beta$ benefits from the FL.
Figure 5: Evaluation of FedSlate versus Baseline Performance under Various Environmental Conditions.
Experimental Results: We conducted multiple iterations of FedSlate and a baseline method under various environmental settings. Specifically, for the parameter selection of $N$ and $n$, we adhered to the standard configurations detailed in the RecSim technical documentation[18]. Our comparison focused on the performance metrics $M_{1}^{{}^{\prime}}$ and $M_{2}^{{}^{\prime}}$. As depicted in Table I, $M_{2}^{{}^{\prime}}$ is consistently less than or equal to $M_{1}^{{}^{\prime}}$ across different settings, indicating that FedSlate can enhance the training velocity of agent $\alpha$. However, it may compromise the agent’s learning of an optimal local policy by potentially reducing the optimal reward. This is attributed to FedSlate’s design, which focuses on optimizing the aggregate long-term benefit for users across platforms rather than maximizing the immediate value for individual users on a single platform. Notably, we omitted the performance comparison between the SlateQ and FedSlate algorithms in the scenario with $N=10,n=3$, since neither algorithm converged in this setting, performing substantially worse than a random recommendation approach. This lack of convergence is likely due to overfitting within an overly simplistic environment. Despite this, We argue that FedSlate effectively aids feedback data owners—designated as agent $\alpha$ in the experiment—by improving training efficiency and significantly reducing computational resource consumption. Moreover, the recommendation strategy learned through FedSlate demonstrates comparable performance to that developed using SlateQ. For a visual representation of the comparative experimental results under different settings, please refer to Fig. 5. In the context of random recommendation algorithms, it is noteworthy that we compare their average reward against the optimal reward achieved by FedSlate and SlateQ, while refraining from providing its ETROR. This approach is inherently justified, given that for random recommendation algorithms, there is no process of ‘learning an optimal recommendation strategy’. Instead, they persistently operate with suboptimal performance, making the utilization of average reward a more representative metric for evaluating the performance of random algorithms.
TABLE I: Comparison of Performance between FedSlate and Baseline under Various Environmental Settings
Metric Method Environment
N=10,n=3 N=100,n=10 N=500,n=10
ETROR SlateQ N/A 3400 2020
FedSlate N/A 2340 1700
Optimal Reward SlateQ N/A 1115.072 1117.974
FedSlate N/A 1106.57 1107.092
Mean Reward Rand 947.087 942.75 938.665
Furthermore, a pivotal aspect of our study is assessing if FedSlate can empower platforms lacking user feedback to augment their recommendation systems by leveraging feedback from other platforms. The performance of agent $\beta$, which employs FedSlate, is contrasted with that of random recommendations. As illustrated in Table II, agent $\beta$ consistently outperforms random recommendations in various settings, as indicated by $R_{M_{3}^{{}^{\prime}}}\geq R_{rnd}$. Please note that our comparative analysis was strictly limited to evaluating the performance of FedSlate-Beta against the random recommendation algorithm. This limitation was necessitated by a fundamental constraint of the original SlateQ framework, which is its incapacity to develop an effective recommendation policy without reward feedback. This finding demonstrates that FedSlate enables entities (represented by agent $\beta$) without feedback data to benefit from user feedback. Such data-deprived entities are likely to have a heightened interest in FedSlate, as it offers a means to exploit insights from feedback data previously inaccessible to them. The comparative results are visually represented in Fig. 6. Notably, in scenarios with $N=10,n=3$, agent $\beta$ demonstrates a reliable learning of the recommendation strategy. This robustness can be attributed to the indirect training process where the target Q-values from agent $\alpha$ differ from agent $\beta$’s objective Q-values, adding a layer of complexity to the training.
Through comparative experiments, we demonstrate the effectiveness of FedSlate for all participating parties in the FL process, indicating that FedSlate can enhance the performance of recommendation systems by leveraging the correlation of user behaviors across different platforms, without compromising user privacy.
TABLE II: Comparison of Performance between FedSlate and Random Recommendation Algorithm under Various Environmental Settings
Environment
Metric Method N=10,n=3 N=100,n=10 N=500,n=10
Optimal Reward FedSlate 1115.281 1102.269 1129.96
Mean Reward Rand 948.294 944.982 939.979
Figure 6: Evaluation of FedSlate versus Random Recommendation Algorithm Performance under Various Environmental Conditions.
(a) Agent Alpha
(b) Agent Beta
Figure 7: Comparison of Performance between Ablated FedSlate Algorithm and Original Algorithm.
(a) Agent Alpha
(b) Agent Beta
Figure 8: Comparison of Performance between Extended FedSlate Algorithm and Original Algorithm.
TABLE III: Comparison of Performance between Ablated FedSlate Algorithm and Original Algorithm
Metric Method Environment
N=100,n=10
ETROR FedSlate-Alpha 2340
FedSlate(abl)-Alpha N/A
FedSlate-Beta 2800
FedSlate(abl)-Beta N/A
Optimal Reward FedSlate-Alpha 1115.072
FedSlate(abl)-Alpha 1005.524
FedSlate-Beta 1102.269
FedSlate(abl)-Beta 956.612
TABLE IV: Comparison of Performance between Extended FedSlate Algorithm and Original Algorithm
Environment(N=100,n=10)
Metric Method dense sparse
FedSlate-Alpha 2340 3540
FedSlate(exp)-Alpha N/A 2280
FedSlate-Beta 2800 2940
ETROR FedSlate(exp)-Beta N/A 780
FedSlate-Alpha 1115.072 1081.422
FedSlate(exp)-Alpha N/A 1091.515
FedSlate-Beta 1102.269 1042.926
Optimal Reward FedSlate(exp)-Beta N/A 1108.855
V-C Ablation Experiment
The FedSlate algorithm employs the extraction of information from the outputs of local networks and its transfer to the global network, which then calculates Q-values used for selecting recommended content. Given the operational mechanism of the algorithm, a concern arises regarding whether the global network tends to discard Q-values that do not originate from itself and solely outputs its own Q-values. Should such a situation occur, it would indicate that our FedSlate algorithm, when making recommendations, solely relies on information from individual platforms, akin to performing SlateQ separately on each platform, without engaging in FL. To investigate this matter, we conduct an ablation experiment.
In the evaluation experiment of FedSlate, both the local and global networks are fully connected networks with five hidden layers. We employ the Mish function [46] as the activation function Nonetheless, the hidden layers vary in size, with the local network possessing a larger size and the global network a smaller size. In the ablation experiment, we maintain the structure of the local network unchanged and simplify the global network as much as possible. We reduce the global network to a single hidden layer network and remove the activation function, transforming it into a simple Linear Model. Through this setup, we force FedSlate to directly utilize Q-values generated by the local network for recommendations. The experimental results, as depicted in Table III, do not provide the ETROR data for the ablated algorithm. Throughout the entire training process, the rewards for the ablation group diverge completely, and its performance does not exhibit significant improvement compared to random recommendation algorithms. The detailed experimental results can be referred to in Fig.7. The ablation experiment demonstrates the effectiveness of our FedSlate framework in terms of vertical FL.
V-D Evaluation of Extended Algorithms
The constructed scenario requires the collaborative efforts of multiple agents to enhance the collective lifetime value of individual users across different platforms. However, in the implementation process, we used the rewards from Platform A to reflect the overall lifetime value. If the team rewards are too sparse, it may adversely affect the performance of FedSlate. To investigate the impact of this factor, we simulated the scenario of sparse team rewards by setting different random seeds for the sub-environments within the main environment, and conducted experiments under the settings of $N=100,n=10$. Additionally, we included the extended version of FedSlate for comparative analysis. The experimental results are presented in Table IV. We observed that when the team rewards in the environment are sparse, the learning process for agents in FedSlate to acquire recommendation policies takes longer, and the resulting policies are less effective. The extended algorithm of FedSlate outperforms the original version in terms of ETROR and Optimal Reward, and we observed that this improvement is more pronounced for agent $\beta$.The visual results of the comparative experiments are shown in Fig. 8. It should be noted that the extended algorithm of FedSlate is merely a supplementary approach for the original algorithm under specific circumstances. It requires agent $\beta$ to have access to reward information, which contradicts the fundamental principles of FedSlate based on vertical FL.
VI Conclusion
To tackle the complexity of integrating user privacy data across diverse platforms into recommendation systems, we introduce a novel reinforcement learning algorithm, designated as FedSlate. This algorithm is designed to develop superior recommendation tactics collaboratively across multiple agents while safeguarding user privacy. Utilizing RecSim, we established a multi-platform recommendation simulation to assess how our algorithm benefits various participants within Federated Learning (FL). Our research indicates that FedSlate effectively resolves the challenges of cross-platform learning in recommendation systems without necessitating the exchange of private data between platforms. Looking ahead, we aim to refine and broaden the scope of FedSlate in the realm of cross-platform recommendation systems that prioritize privacy. To this end, we plan to augment FedSlate with cutting-edge privacy protection methods, such as secure aggregation, differential privacy, and FL with encrypted data, as detailed by Lyu et al. (2022)[47]. These enhancements will ensure the integrity of user privacy throughout the collaborative learning process, even when handling sensitive data. Moreover, we will undertake comprehensive experiments and evaluations of the FedSlate algorithm in authentic recommendation settings involving diverse platforms and user bases. This will yield critical insights into the algorithm’s efficiency, scalability, and practical applicability, alongside its influence on user satisfaction and engagement.
VII Acknowledgement
This work is supported by the National Natural Science Foundation of China, The Research Project of Shanghai Science and Technology Commission (Grant No.62102241, No.23ZR1425400).